# Local Look-Ahead Guidance via Verifier-in-the-Loop for Automated Theorem Proving

### v1

@@ Abstract
The most promising recent methods for AI reasoning require applying variants of reinforcement learning (RL) either on rolled out trajectories from the model, even for the step-wise rewards, or large quantities of human annotated trajectory data. The reliance on the rolled-out trajectory renders the compute cost and time prohibitively high. In particular, the correctness of a reasoning trajectory can typically only be judged at its completion, leading to sparse rewards in RL or requiring expensive synthetic data generation in expert iteration-like methods. In this work, we focus on the Automatic Theorem Proving (ATP) task and propose a novel verifier-in-the-loop design, which unlike existing approaches that leverage feedback on the entire reasoning trajectory, employs an automated verifier to give intermediate feedback at each step of the reasoning process. Using Lean as the verifier, we empirically show that the step-by-step local verification produces a global improvement in the model’s reasoning accuracy and efficiency.
@@ 1 Introduction
As the new applications of modern machine learning are emerging in various scientific and engineering domains, automated mathematical theorem proving has garnered interests from both machine learning researchers and mathematicians. Many ongoing efforts leverage reinforcement learning and expert iteration, inspired by the success of methods like AlphaZero, to build models that search the proof space and provide step wise or holistic solutions (Lample et al., 2022; Xin et al., 2024a; Gloeckle et al., 2024; Anthony et al., 2017; Silver et al., 2018). These solutions are usually verified by formal proof verification systems like Lean (Moura and Ullrich, 2021) or Coq (Coq Development Team, 2024). Relying on Reinforcement Learning (RL) is advantageous in terms of data efficiency but comes with high computational and training costs (Gloeckle et al., 2024). Part of this complexity is related to the necessity of rolling out the proofs and computing rewards from successful episodes.
@@
In contrast, ReProver (Yang et al., 2023) takes a simpler supervised training approach, specifically imitation learning, paired with premise retrieval methods. The key components of this proof system are as follows: theorem that we would like to prove, tactics that are actions toward the final proof and itself consist of set of goals to be proven, premises that are used to prove goals, and state of the proof which includes the set of goals that are still unproven. The approach consists of retrieving the relevant premises from a database given the final theorem and the state of the proof, and then using ReProver to provide tactics for getting to the next state. The proof terminates when all the goals are proven. The method achieves competitive performance with an order of magnitude smaller complexity and training time (Gloeckle et al., 2024). While the computational cost and simplicity of ReProver are appealing, we empirically observed that many failure cases of ReProver are due to syntactically incorrect tactics or tactics that are not applicable to the current state of a proof. This has detrimental effects on the beam search performed at inference time, as many beams result in invalid tactics that need to be verified by Lean, thus taking away time from exploring more promising tactics.
@@
To preserve the desirable computational efficiency of ReProver and simultaneously address this problem, the natural choice is to fine-tune the model to remove syntactic errors and increase the number of useful tactics for the proof at each state. Recently, feedback-based alignment has been gaining traction in various other similar fields such as automated code generation and various preference optimization methods where the rewards come either from human feedback (RLHF) or other reward models (Ouyang et al., 2022; Ziegler et al., 2019; Rafailov et al., 2024). Since applying many such reinforcement learning methods for training large models can be complex and expensive, various methods have been introduced in the literature with moderate complexity, among which we can refer to Direct Preference Optimization (DPO(Rafailov et al., 2024) and Group Reward Preference Optimization (GRPO(Shao et al., 2024). In context of mathematical reasoning and theorem proving, many works emphasized the importance of trajectory-level preferences in mathematical problem solving with large language models (LLMs) (e.g. see Xiong et al. (2024) and Preference Optimization paragraph in Sec. 2). Even when the complexity is saved in the RL training algorithm, computing these trajectory level preferences can incur additional complexity. This discussion extends to more general episodic reasoning tasks with stepwise verification where the model needs to provide outputs that are both syntactically and semantically correct and useful for solving the problem at hand.
@@
In this work, we aim to address the above issues by fine-tuning a pre-trained model which listens and uses the feedback from the tool, in this case Lean, during the training. At each step, our framework, called LeanListener, obtains feedback on the generated tactics directly via its interaction with the Lean software, and performs policy optimization with a reward that is designed based on Lean feedback. Given the pre-trained ReProver model from Yang et al. (2023), we sample different tactics from its output for each proof state in the training set and use Lean feedback to compute the reward. The reward consists of a negative return for invalid tactics, a positive one for applicable tactics, and a return based on the number of remaining unsolved goals. The RL training is done using GRPO. First note that the sampling step for ReProver’s output can yield applicable and new tactics that differ from the provided human label. Therefore, it helps the data efficiency of the method by exploring and adding new tactics like what we see in expert iteration. Second, unlike methods like Process-supervised Reward Model (PRM), we do not compute the step-wise reward based on the full trajectory information and only rely on local look-ahead feedback from the number of remaining and unsolved goals in the next immediate steps, and therefore, we address the complexity of the trajectory based preference association. Thanks to this fine-tuning strategy, we expect the model to rank valid and effective tactics higher than invalid ones, even if they were not previously observed in the human-labeled trajectories. As a result, our fine-tuned model can make better use of the beam search used at inference time. Besides, as we will show in our numerical results, the local look ahead, online training with Lean in loop, and GRPO are crucial components in improving the final performance of the model. To summarize, our contributions are as follows. We propose a framework for efficient training of the ReProver to leverage the feedback provided by the external tool, in this case, the Lean solver. We use GRPO for training the model using the reward that is based on the applicable tactics and the number of unsolved goals. The GRPO training is particularly beneficial compared to DPO training in our case. We show that the training based on offline dataset generation of positive-negative samples is ineffective, and online training is crucial. The online verification from the tool automatically adds new tactics to the training data. Finally, we can obtain noticeable gains by computing the reward using local look-ahead without recourse to the trajectory level preference.
@@
Figure 1: We start with a dataset of human-annotated theorems and their proofs, including the gold tactics, in Lean. During training, the ReProver retrieves the relevant premises from the premise database, and based on those and the state of the proof generates the next tactics. The generated tactics are verified by Lean, and its feedback is used by RL training loop to fine-tune the ReProver. The loop continues using the steps from the human annotated tactics until proof termination.
@@ 2 Related Works
Machine learning is currently used in automatic theorem proving to generate the whole proof, or provide assistance in sequential proofs by finding the right premise or right tactics, or assisting in theorem prover specific formalization of mathematical statements, as well as combination of these ideas; see Xin et al. (2024a; b); Polu and Sutskever (2020); Jiang et al. (2022a); Wang et al. (2023); Jiang et al. (2022b); Wu et al. (2022) for some pointers. We review some of these works in more details. A key component for theorem proving is the formal proof management systems and verifiers like Lean Moura and Ullrich (2021) and Coq Coq Development Team (2024). Lean has received particular attention in the mathematics community for example by its use in verification of some components in the condensed mathematics research program Scholze (2022). We use Lean in this work, and a framework that consists of a premise retrieval part and a prover that generates step-wise proof tactics. In our paper, we are focused on the logic behind tactic generation. In this sense, our approach can be complementary to any work on retrieval and search steps.
Searching over different proofs and tactics is another component explored in various works, for example Polu et al. (2023); Xin et al. (2024b); Lample et al. (2022). As discussed in Polu et al. (2023), any RL algorithm for theorem proving should address two challenges, namely an infinite dimensional action space, and the absence of a natural opponent for self-play. Therefore, the authors suggest using expert iteration Anthony et al. (2017), which amounts to iteratively fine-tuning a based model and searching to generate correct proofs. They also introduce lean-gym to facilitate the search procedure in Lean. HyperTree Proof Searchs was introduced in Lample et al. (2022) which is a new search algorithms within an online reinforcement learning procedure. The key idea is that the search is represented over a hypergraph, where a policy network generates tactics composed of sub-goals, and then each goal is expanded for proof using new tactics. The provability of each goal is approximated using a critic. The idea of keeping the visit counts, action values and their statistics follows the similar MCTS procedure. In Gloeckle et al. (2024), the authors introduce a reinforcement learning based framework for theorem proving in Lean 4, which consist of using a programming interface based on Aesop for proof search organization, the HTPS (Lample et al. (2022)) procedure with an online reinforcement learning step. In Zhao et al. (2023), the authors use subgoal learning from reinforcement learning to decompose an informal LLM generated proof into subgoals and verify using a verifier. They also use a diffusion model for demonstration organization.
The transformer-based LLMs are used also for theorem proving. For example GPT-f was introduced in Polu and Sutskever (2020) using Metamath as the formalization framework. They showed iterative training a value function on proofs generated by the prover can continually improve the performance. Another example is LLEMMA, which is based on pretrained Llama Code Azerbayev et al. (2024).
@@
DeepSeek-Prover Xin et al. (2024a) generated Lean4 proof data by translating natural language problems into their formal versions in Lean4, and the model produces whole proofs in single turns. DeepSeekV1.5 Xin et al. (2024b) provides a middle ground by generating the whole proof, verifying it via Lean, and then truncating the proof until the first error. The authors propose additional techniques relying on proper appending of previous states and including truncate-and-resume in the Monte-Carlo Tree Search (MCTS) procedure. Additionally, DeepSeekV1.5 leverages verification feedback from Lean on whole-proofs to improve model’s alignment with the formal structure of Lean. In particular, it uses the Group-Relative Policy Optimization (GRPO) algorithm originally introduced for DeepSeek-Math in Shao et al. (2024), which removes the need for a separate critic model, thereby simplifying the reinforcement learning pipeline and reducing its computational cost. In comparison, we consider a step-wise approach, leveraging the verifier feedback at each intermediate step of the proof.
@@
Other works in the literature have considered the idea of model alignment to improve the reasoning capabilities of language models, too. In general, there are online version of Reinforcement Learning with Human Feedback (RLHF) as in Bai et al. (2022); Ouyang et al. (2022), and offline versions with either an explicit reward model (e.g. Christiano et al. (2017); Ziegler et al. (2019)) or an implicit reward model (e.g. DPO Rafailov et al. (2024)) to encode the preference. The later method directly optimizes the model without training any independent reward model. In particular DPO uses Bradley-Terry preference model Bradley and Terry (1952) and the fact that the optimal solution to the KL-constrained reward maximization objective is known in closed form to simplify the training loss without dependence on a reward model Rafailov et al. (2024). Many follow-up works explored DPO variations and its shortcomings such as reward over-optimization, for example by introducing Kahneman-Tversky Optimization (KTO) Ethayarajh et al. (2024) and introducing trust-region based versions of previous alignment models Gorbatovski et al. (2024).
@@
3 LeanListener
@@
In Interactive Theorem Proving (ITP), an LLM-based prover interacts with an external proof assistant and receives feedback on the steps to be taken to prove the given theorem. In this work, we use the proof assistant software Lean [^1] We use Lean 4, and for brevity, simply refer to it as Lean in the rest of the manuscript. (de Moura et al., 2015; Moura and Ullrich, 2021) as the formal environment which LLMs employ to verify each proof step in a proof sequence. In particular, we used the open-source LeanDojo framework (Yang et al., 2023), which provides toolkits to interact with Lean, as well as readily extracted data and pre-trained models.
@@
We model the theorem proving task as a Markov Decision Process (MDP), a tuple $(S,A,P,R)$, in which $S$ is the set of proof states, $A$ is the set of all possible tactics (i.e., proof steps), $P$ is the transition probability function that describes the likelihood of moving from one state to another given a specific tactic, and $R$ is the reward function that assigns a value to each state transition. An agent interacts with Lean to prove a theorem $T$ by iteratively observing the current ”proof state” $s_{t}$ and performing an action $a_{t}$, known as a ”tactic”. This action transitions the proof to the next state $s_{t+1}$. Lean evaluates each tactic, either accepting or rejecting it, and provides additional feedback, such as the number of goals remaining to be proven. Our model leverages this feedback during training and for generating subsequent tactics. It is important to note that Lean is a formal language with strict syntax rules, meaning only certain tactics are applicable in each state. At each step, the agent relies solely on Lean’s feedback about the current state to generate the next tactic until the theorem $T$ is proven or the model exhausts the predefined time limit. Figure 2 presents a Lean proof example. Finally, the sequence of state-tactic pairs $\{\langle s_{t},a_{t}\rangle\}_{t}$ composes the proof of the original theorem $T$.
@@
In general, a certain state $s_{t}$ can include multiple (sub-) goals, representing a number of independent statements to be simultaneously proved in order to prove the original statement. An intuitive example is given by application of an induction tactic, which turns a single statement into two independent ones (the base case and the inductive step). Hence, the application of a tactic $a_{t}$ on a certain state $s_{t}$ can prove certain goals, but also turn some into more sub-goals. As a result, the number of sub-goals $\mathcal{G}(s_{t+1})$ can increase, decrease or remain unchanged after the application of each tactic.
@@
Tactic prediction is one of the primary tasks in training LLM provers (Yang et al., 2023; Welleck and Saha, 2023; Lample et al., 2022). In a supervised learning framework, during training, the model is fed the current state ($s_{t}$) and asked to predict a tactic ($a_{t}$). Recent studies have shown that providing related premises, in addition to the current state, can also enhance the model’s performance (Mikuła et al., 2023; Yang et al., 2023). In the inference phase, however, for each theorem, the model employs an inference time compute technique. It observes the current state ($s_{t}$) and generates a set of tactics ($\set{a_{0},...,a_{k}}$) using beam search with a size of $k$. Building a proof search tree, the tactic generator interacts with the Lean assistant, starting with the tactic with the highest accumulative log probability (best first search). The returned state can be an error state if the tactic execution is unsuccessful, e.g., due to timeout or inapplicable tactic. In this case, the model explores the next tactic in the queue. This process continues until we reach a valid next state ($s_{t+1}$) or prove the given theorem. It is worth mentioning that the generated proof by the model can be different from the human-written one in the dataset. As many inference-time compute approaches, the model (generator) utilizes feedback from an external expert (verifier) and explores creative proofs at cost of extra computation at inference.
@@
While the pre-training objective is a supervised sequence-to-sequence task, in the evaluation phase, the model acts like an RL agent, which receives feedback on its actions (generated tactics) and uses it to choose the next action via the search strategy explained above. In this work, we resolve this discrepancy and bring the proof-tree expansion based exploration to the training regime, leveraging the feedback provided to each tactic to construct step-wise rewards. Inspired by human preference in LLMs alignment, we present a novel framework, LeanListener, in which we align the pre-trained model with guidance from Lean in a step-wise scenario. In contrast with similar works with RL training like Xin et al. (2024b), we only need a single step look ahead, and do not need the full trajectory to compute the rewards. In what follows, we explain our LeanListener framework in detail.
Figure 2: On the left, a proof example in Lean shows the $min$ function is symmetric, and on the right, its proof tree. The numbers next to the boxes refer to corresponding line numbers in the written proof.
3.1 Methodology

### v2

@@ Abstract
The most promising recent methods for AI reasoning require applying variants of reinforcement learning (RL) either on rolled out trajectories from the LLMs, even for the step-wise rewards, or large quantities of human-annotated trajectory data. The reliance on the rolled-out trajectory renders the compute cost and time prohibitively high. In particular, the correctness of a reasoning trajectory can typically only be judged at its completion, leading to sparse rewards in RL or requiring expensive synthetic data generation in expert iteration-like methods. In this work, we focus on the Automatic Theorem Proving (ATP) task and propose a novel verifier-in-the-loop design, which, unlike existing approaches that leverage feedback on the entire reasoning trajectory, employs an automated verifier to give intermediate feedback at each step of the reasoning process. Using Lean as the verifier, we empirically show that the step-by-step local verification produces a global improvement in the model’s reasoning accuracy and efficiency.
Local Look-Ahead Guidance via Verifier-in-the-Loop for Automated Theorem Proving
Sara Rajaeethanks: Equal contribution.thanks: Work done during an internship at Qualcomm AI Research, Amsterdam. 1, Kumar Pratik2, Gabriele Cesa2, Arash Behboodi2 Language Technology Lab, University of Amsterdam1 Qualcomm AI Researchthanks: Qualcomm AI Research is an initiative of Qualcomm Technologies, Inc.    ©2025 Qualcomm Technologies, Inc. and/or its affiliated companies. All Rights Reserved.  , Amsterdam2 s.rajaee@uva.nl {kpratik, gcesa, behboodi}@qti.qualcomm.com
@@ 1 Introduction
As the new applications of modern large language models (LLMs) are emerging in various scientific and engineering domains, automated mathematical theorem proving has garnered interest from both machine learning researchers and mathematicians. Many ongoing efforts leverage reinforcement learning and expert iteration, inspired by the success of methods like AlphaZero, to build models that search the proof space and provide step-wise or holistic solutions Lample et al. (2022); Xin et al. (2024a); Gloeckle et al. (2024); Anthony et al. (2017); Silver et al. (2018). These solutions are usually verified by formal proof verification systems like Lean Moura and Ullrich (2021) or Coq Coq Development Team (2024). Relying on Reinforcement Learning (RL) is advantageous in terms of data efficiency but comes with high computational and training costs Gloeckle et al. (2024). Part of this complexity is related to the necessity of rolling out the proofs and computing rewards from successful episodes.
@@
In contrast, ReProver Yang et al. (2023) takes a simpler supervised training approach, specifically imitation learning, paired with premise retrieval methods. The key components of this proof system are as follows: the theorem we would like to prove, tactics that are actions toward the final proof and itself consist of a set of goals to be proven, premises that are used to prove goals, and the state of the proof which includes the set of goals that are still unproven. The approach consists of retrieving the relevant premises from a database given the final theorem and the state of the proof, and then using ReProver to provide tactics for getting to the next state. The proof terminates when all the goals are proven. The method achieves competitive performance with an order of magnitude smaller complexity and training time Gloeckle et al. (2024). While the computational cost and simplicity of ReProver are appealing, we empirically observed that many failure cases of tactic generators like ReProver are due to syntactically incorrect tactics or tactics that are not applicable to the current state of a proof. This has detrimental effects on the beam search performed at inference time, as many beams result in invalid tactics that need to be verified by Lean, thus taking away time from exploring more promising tactics.
@@
To preserve the desirable computational efficiency of the ReProver and simultaneously address this problem, the natural choice is to fine-tune the model to remove syntactic errors and increase the number of useful tactics for the proof at each state. Recently, feedback-based alignment has been gaining attraction in various other similar fields, such as automated code generation and various preference optimization methods where the rewards come either from human feedback (RLHF) or other reward models Ouyang et al. (2022); Ziegler et al. (2019); Rafailov et al. (2024). Since applying many such reinforcement learning methods for training large models can be complex and expensive, various approaches have been introduced in the literature with moderate complexity, among which we can refer to Direct Preference Optimization (DPORafailov et al. (2024) and Group Reward Preference Optimization (GRPOShao et al. (2024). In the context of mathematical reasoning and theorem proving, many works have emphasized the importance of trajectory-level preferences in mathematical problem solving with large language models (LLMs) (e.g., see Xiong et al. (2024) and Preference Optimization paragraph in Sec. 2). Even when complexity is saved in the RL training algorithms, computing these trajectory-level preferences can incur additional complexity. This discussion extends to more general episodic reasoning tasks with stepwise verification, where the model needs to provide outputs that are both syntactically and semantically correct and useful for solving the problem at hand.
@@
In this work, we aim to address the above issues by fine-tuning a pre-trained prover, which listens and uses the feedback from the tool, in this case Lean, during the training. At each step, our framework, called LeanListener, obtains feedback on the generated tactics directly via its interaction with the Lean software, and performs policy optimization with a reward that is designed based on Lean feedback. Given the pre-trained tactic generator from Yang et al. (2023), we sample different tactics from its output for each proof state in the training set and use Lean feedback to compute the reward. The reward consists of a negative return for invalid tactics, a positive one for applicable tactics, and a return based on the number of remaining unsolved goals. The RL training is done using GRPO. First, note that the sampling step for the generator’s output can yield applicable and new tactics that differ from the provided human label. Therefore, it helps the data efficiency of the method by exploring and adding new tactics like what we see in expert iteration. Second, unlike methods like Process-supervised Reward Model (PRM), we do not compute the step-wise reward based on the full trajectory information and only rely on local look-ahead feedback from the number of remaining and unsolved goals in the next immediate steps and, therefore, we address the complexity of the trajectory based preference association. Thanks to this fine-tuning strategy, we expect the model to rank valid and effective tactics higher than invalid ones, even if they were not previously observed in the human-labeled trajectories. As a result, our fine-tuned model can make better use of the beam search used at inference time. Besides, as we will show in our numerical results, the local look ahead, online training with Lean-in-loop, and GRPO are crucial components in improving the final performance of the model. To summarize, our contributions are as follows. We propose a framework for efficient training of a tactic generator LLM to leverage the feedback provided by the external tool, in this case, the Lean solver. We use GRPO for training the model using the reward that is based on the applicable tactics and the number of unsolved goals. The GRPO training is particularly beneficial compared to DPO training in our case. We show that the training based on offline dataset generation of positive-negative samples is ineffective, and online training is crucial. The online verification from the tool automatically adds new tactics to the training data. Finally, we can obtain noticeable gains by computing the reward using local look-ahead without recourse to the trajectory level preference.
@@
Figure 1: We start with a dataset of human-annotated theorems and their proofs, including the gold tactics, in Lean. During training, the ReProver retrieves the relevant premises from the premise database, and based on those and the state of the proof, generates the next tactics. The generated tactics are verified by Lean, and its feedback is used by RL training loop to fine-tune the ReProver. The loop continues using the steps from the human-annotated tactics until proof termination.
@@ 2 Related Works
Large Language Models (LLMs) are currently used in automatic theorem proving to generate the whole proof, or provide assistance in sequential proofs by finding the right premise or right tactics, or assisting in theorem prover specific formalization of mathematical statements, as well as combination of these ideas; see Xin et al. (2024a, b); Polu and Sutskever (2020); Jiang et al. (2022a); Wang et al. (2023); Jiang et al. (2022b); Wu et al. (2022) for some pointers. A key component for theorem proving is the formal proof management systems and verifiers like Lean Moura and Ullrich (2021) and Coq Coq Development Team (2024).
@@
In Thor Jiang et al. (2022a), the authors introduce a class of methods to use automated theorem provers for premise selection. LeanDojo Yang et al. (2023) provided a framework that retrieves the premise from a database of Lean premises and uses ReProver, an encoder-decoder model, to generate tactics for theorem proving in Lean. LeanAgent Kumarappan et al. (2024) adds a dynamic database to LeanDojo, which enables the continual learning of the agent. Besides, a new curriculum learning based on the difficulty of the theorems is also added. GPT-f was introduced in Polu and Sutskever (2020) using Metamath as the formalization framework. Besides, they showed that iterative training a value function on proofs generated by the LLM prover can continually improve the performance. Another example is LLEMMA, which is based on pre-trained Llama Code Azerbayev et al. (2024). DeepSeek-Prover Xin et al. (2024a) generated Lean4 proof data by translating natural language problems into their formal versions in Lean4, and the model produces whole proofs in single turns. DeepSeekV1.5 Xin et al. (2024b) provides a middle ground by generating the whole proof, verifying it via Lean, and then truncating the proof until the first error. The authors propose additional techniques relying on proper appending of previous states and including truncate-and-resume in the Monte-Carlo Tree Search (MCTS) procedure. Additionally, DeepSeekV1.5 leverages verification feedback from Lean on whole-proofs to improve model’s alignment with the formal structure of Lean. In particular, it uses the Group-Relative Policy Optimization (GRPO) algorithm originally introduced for DeepSeek-Math in Shao et al. (2024), which removes the need for a separate critic model, thereby simplifying the reinforcement learning pipeline and reducing its computational cost. In comparison, we consider a step-wise approach, leveraging the verifier feedback at each intermediate step of the proof.
@@
Other works in the literature have considered the idea of model alignment to improve the reasoning capabilities of language models, too. In general, there are online versions of Reinforcement Learning with Human Feedback (RLHF) as in Bai et al. (2022); Ouyang et al. (2022), and offline versions with either an explicit reward model (e.g. Christiano et al. (2017); Ziegler et al. (2019)) or an implicit reward model (e.g. DPO Rafailov et al. (2024)) to encode the preference. The later method directly optimizes the model without training any independent reward model. Many follow-up works explored DPO variations and its shortcomings, such as reward over-optimization, for example, by introducing Kahneman-Tversky Optimization (KTO) Ethayarajh et al. (2024).
@@
3 Methodology
In this section, we begin with an overview of Interactive Theorem Proving (ITP), outlining the task formulation and key terminology. We then introduce our framework, LeanListener, and describe its main components.
3.1 Background
@@
In ITP, an LLM-based prover interacts with an external proof assistant and receives feedback on the steps to be taken to prove the given theorem. In this work, we use the proof assistant software Lean [^1] We use Lean 4, and for brevity, simply refer to it as Lean in the rest of the manuscript. de Moura et al. (2015); Moura and Ullrich (2021) as the formal environment which LLMs employ to verify each proof step in a proof sequence. In particular, we used the open-source LeanDojo framework Yang et al. (2023), which provides toolkits to interact with Lean, as well as readily extracted data and pre-trained models.
@@
An agent can interact with Lean to prove a theorem $T$ by iteratively observing the current "proof state" $s_{t}$ and performing an action $a_{t}$, called a "tactic" (i.e., proof steps), which leads to the next state $s_{t+1}$. Lean inspects each tactic, accepts or rejects it, and provides additional information, like the number of goals yet to be proven. Our model can utilize the feedback received from Lean during the training and for the next tactic generation. Note that Lean is a formal language; hence, it has a strict syntax, and only certain tactics can be applied on each state. At each step, the agent relies only on the Lean feedback about the current state to generate the next tactic until the theorem $T$ is proved or the model runs out of predefined time. Figure 5 presents a Lean proof example. Finally, the sequence of state-tactic pairs $\{\langle s_{t},a_{t}\rangle\}_{t}$ composes the proof of the original theorem $T$.
@@
In general, a certain state $s_{t}$ can include multiple (sub-) goals, representing a number of independent statements to be simultaneously proved in order to prove the original statement. An intuitive example is given by application of an induction tactic, which turns a single statement into two independent ones (the base case and the inductive step). Hence, the application of a tactic $a_{t}$ on a certain state $s_{t}$ can prove certain goals but also turn some into more sub-goals. As a result, the number of sub-goals $\mathcal{G}(s_{t+1})$ can increase, decrease or remain unchanged after the application of each tactic.
@@
Tactic prediction is one of the primary tasks in training LLM provers Yang et al. (2023); Welleck and Saha (2023); Lample et al. (2022). In a supervised learning framework, during training, the model is fed the current state ($s_{t}$) and asked to predict a tactic ($a_{t}$). Recent studies have shown that providing related premises, in addition to the current state, can also enhance the model’s performance Mikuła et al. (2023); Yang et al. (2023). In the inference phase, however, for each theorem, the model employs an inference time-compute technique. It observes the current state ($s_{t}$) and generates a set of tactics ($\set{a_{0},...,a_{k}}$) using beam search with a size of $k$. Building a proof search tree, the tactic generator interacts with the Lean assistant, starting with the tactic with the highest accumulative log probability (best first search). The returned state can be an error state if the tactic execution is unsuccessful, e.g., due to timeout or inapplicable tactic. In this case, the model explores the next tactic in the queue. This process will continue until we reach a valid next state ($s_{t+1}$) or prove the given theorem. It is worth mentioning that the generated proof by the model can be different from the human-written one in the test dataset. As many inference-time compute approaches, the model (generator) utilizes feedback from an external expert (verifier) and can explore creative proofs at the cost of extra computation at inference.
While the pre-training objective is a supervised sequence-to-sequence task, in the evaluation phase, the model acts like an RL agent, which receives feedback on its actions (generated tactics) and uses it to choose the next action via the search strategy explained above.
@@
In this work, we resolve this discrepancy and bring the proof-tree expansion-based exploration to the training regime, leveraging the feedback provided to each tactic to construct step-wise rewards. Inspired by human preference in LLMs alignment, we present a novel framework, LeanListener, in which we align the pre-trained tactic generator with guidance from Lean in a step-wise scenario. In contrast with similar works with RL training like Xin et al. (2024b), we only need a single step look ahead and do not utilize the full trajectory to compute the rewards. In what follows, we explain our LeanListener framework in detail.
3.2 LeanListener